# DAFE: LLM-Based Evaluation Through Dynamic Arbitration for Free-Form Question-Answering

### v1

@@ C.8 DAFE in multi-reference answers
C.9 Analysis of automatic metrics
Figures 8, 9, 10, and 11 illustrate the fundamental trade-offs in automatic metrics. In TriviaQA, where multiple normalized reference answers exist, EM achieves impressive true positives (61.7-74.3%) compared to HotpotQA (23.0-34.3%) which contains single reference answers. EM’s near-zero false positives across tasks (0-0.7%) stem from its strict string matching – it only flags matches when answers are identical to references. Our error analysis found three primary causes of such rare false positives including preprocessing errors, where character normalization removes crucial distinctions, and reference ambiguities, where incomplete or ambiguous references lead to incorrect matches. Additionally, a semantic mismatch occurs when the EM incorrectly labels a prediction as true by matching text without considering its context. For instance, despite their different contextual meanings, EM wrongly marks a match between a model prediction of “1944” (describing the start of a war) and a reference answer containing “1944” (representing the end of the war).
@@ C.9 Analysis of automatic metrics
Appendix D Prompting
In our main experiment, we performed zero-shot prompting in the following two stages.
D.1 Prompting Candidate LLMs
We prompted candidate LLMs (see Figure 12) to record generations for each task. We set the same role and prompt structure for each candidate model to ensure the reproducibility of our results. Figure 13 shows the candidate GPT-3.5-turbo response at zero temperature for the input given in Figure 12.
You are a helpful assistant. Which comic book was also written by the writer of Crossed?
Figure 12: Prompting candidate GPT-3.5-turbo to elicit outputs for HotpotQA.
Crossed was written by Garth Ennis. The Boys is a comic book series also written by Garth Ennis. So the comic book that was also written by the writer of Crossed is The Boys.
Figure 13: Candidate GPT-3.5-turbo response to the input given in Figure 12.
D.2 Prompting LLM Judges
We prompted LLMs-as-judges to perform the evaluation (see Figure 14). In Figure 15, judge Llama-3.1-70B evaluating candidate GPT-3.5-turbo.
You are a helpful assistant acting as an impartial judge. You will be given a Question and a Proposed Answer. Your task is to judge whether the Proposed Answer is correct by comparing it to the Reference Answer. If the Proposed Answer is correct, choose ’True’, otherwise choose ’False’. Provide a brief explanation for your decision. Question: Which comic book was also written by the writer of Crossed? Provided Answer: Crossed was written by Garth Ennis. The Boys is a comic book series also written by Garth Ennis. So the comic book that was also written by the writer of Crossed is The Boys. So the answer is: The Boys. Reference Answer: the Vertigo series “Preacher” Evaluation: Provide your response in the following format: Decision: [True/False] Explanation: [Your brief explanation]
Figure 14: Prompting judge Llama-3.1-70B for evaluating candidate GPT-3.5-turbo.
Decision: False Explanation: The Proposed Answer incorrectly identifies "The Boys" as the comic book written by the writer of Crossed. While it is true that Garth Ennis wrote Crossed, the Proposed Answer fails to consider other works written by Garth Ennis. The Reference Answer, "Preacher", is indeed another comic book series written by Garth Ennis, specifically for the Vertigo imprint.
Figure 15: Llama-3.1-70B-Judge verdict on the candidate GPT-3.5-turbo output.

### v2

@@ E.7 CLEV in multi-reference answers
E.8 Analysis of automatic metrics
Figures 3, 4, 5, and 6 illustrate the fundamental trade-offs in automatic metrics. In TriviaQA, where multiple normalized reference answers exist, EM achieves impressive true positives (61.7-74.3%) compared to HotpotQA (23.0-34.3%) which contains single reference answers. EM’s near-zero false positives across tasks (0-0.7%) stem from its strict string matching – it only flags matches when answers are identical to references. Our error analysis found three primary causes of such rare false positives including preprocessing errors, where character normalization removes crucial distinctions, and reference ambiguities, where incomplete or ambiguous references lead to incorrect matches. Additionally, a semantic mismatch occurs when the EM incorrectly labels a prediction as true by matching text without considering its context. For instance, despite their different contextual meanings, EM wrongly marks a match between a model prediction of “1944” (describing the start of a war) and a reference answer containing “1944” (representing the end of the war).
@@ E.8 Analysis of automatic metrics
Appendix F Prompting
In our main experiment, we performed zero-shot prompting in the two stages. First, we prompted candidate LLMs to obtain outputs for the given asks. Then, we instructed LLM judges to perform evaluations.
Candidate Prompt You are a helpful assistant. Which comic book was also written by the writer of Crossed?
Figure 7: Prompt given to candidate GPT-3.5-turbo to elicit an answer for HotpotQA.
Candidate Response Crossed was written by Garth Ennis. The Boys is a comic book series also written by Garth Ennis. So the comic book that was also written by the writer of Crossed is The Boys.
Figure 8: Candidate GPT-3.5-turbo response to the prompt shown in Figure 7.
Judge Prompt You are a helpful assistant acting as an impartial judge. You will be given a Question and a Proposed Answer. Your task is to judge whether the Proposed Answer is correct by comparing it to the Reference Answer. If the Proposed Answer is correct, choose ’True’, otherwise, choose ’False’. Provide a brief explanation for your decision. Question: Which comic book was also written by the writer of Crossed? Provided Answer: Crossed was written by Garth Ennis. The Boys is a comic book series also written by Garth Ennis. So the comic book that was also written by the writer of Crossed is The Boys. Reference Answer: the Vertigo series “Preacher” Evaluation: Provide your response in the following format: Decision: [True/False] Explanation: [Your brief explanation]
Figure 9: Prompting judge Llama-3.1-70B for evaluating the candidate GPT-3.5-turbo answer.
Judge Response Decision: False Explanation: The Proposed Answer incorrectly identifies ”The Boys” as the comic book written by the writer of Crossed. While it is true that Garth Ennis wrote Crossed, the Proposed Answer fails to consider other works written by Garth Ennis. The Reference Answer, ”Preacher”, is indeed another comic book series written by Garth Ennis, specifically for the Vertigo imprint.
Figure 10: Llama-3.1-70B-Judge verdict on the candidate GPT-3.5-turbo answer.